# Direct reconstruction of anisotropic self-adjoint inclusions in the Calderón problem

### v1

@@ 6. Localized potentials for anisotropic coefficients
$\int_{\Omega}A^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V}F\cdot\overline{\nabla v}$ (6.1)
@@
We let $V_{1}\subseteq B\subset V\subseteq U$ where $V_{1}$ and $V$ are relatively open non-empty connected sets, and define $V_{2}=\Omega\setminus\overline{V}$. These sets are chosen such that $V$ intersects $\Gamma$, $\mathop{\textup{dist}}(V_{1},V_{2})>0$, and that $\partial(\Omega\setminus\overline{V_{j}})$ is Lipschitz continuous.
We start by proving that $R(L_{V_{1}}(A))$ and $R(L_{V_{2}}(A))$ have a trivial intersection. Hence, let $\varphi\in R(L_{V_{1}}(A))\cap R(L_{V_{2}}(A))$, i.e. $\varphi=w_{j}|_{\Gamma}$ with $w_{j}=w_{F_{j}}^{A}$ for some $F_{j}\in L^{2}(V_{j})^{d}$ for $j\in\{1,2\}$. From (6.1) we conclude that
$\int_{\Omega}A^{*}\nabla w_{j}\cdot\overline{\nabla v}\,\mathrm{d}x=0\text{ for all }v\in H^{1}(\Omega\setminus\overline{V_{j}}).$
From the weak form of the conductivity problem (2.2), with conductivity $A^{*}$, we thus have
@@
$\nu\cdot(A^{*}\nabla w_{j})$ $=0\text{ on }\partial(\Omega\setminus\overline{V_{j}}),$
@@
The following result shows that the sequence $(f_{j})$ of Neumann conditions (e.g. from Theorem 6.3) such that the corresponding electric power $\lvert\nabla u_{j}\rvert^{2}$ is localized, can be chosen independently of the function values of $A$ on the set where the power tends to zero.
@@
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset D^{\bullet}\cap U$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
@@
Case 1: Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{+}_{C}}$ satisfy
@@
Case 2: Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{-}_{C}}$ satisfy
@@
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset D^{\bullet}\cap U$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
@@
Case 1: We now combine Proposition 5.5 with Proposition 5.6(ii), using the facts that $A_{D}-A_{0}$ is positive semidefinite in $U$ and $A_{D}-A_{0}\geq cI$ in $B$ for some scalar $c>0$, which results in
@@
Case 2: We use Proposition 5.5 together with the facts that $A_{D}-A_{0}$ is negative semidefinite in $U$ and $A_{D}-A_{0}\leq-cI$ in $B$ for some scalar $c>0$, which results in
@@
Next, assume $B\not\subseteq D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
@@
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{implies}\quad B\subseteq D^{\bullet}.$
@@
Next, assume $B\not\subseteq D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy

### v2

@@ 6. Localized potentials for anisotropic coefficients
$\int_{\Omega}A^{*}\nabla w\cdot\overline{\nabla v}\,\mathrm{d}x=\int_{V}F\cdot\overline{\nabla v}\,\mathrm{d}x$ (6.1)
@@
We let $V_{1}\Subset B\subset V\subseteq U$ where $V_{1}$ and $V$ are relatively open non-empty connected sets, and define $V_{2}=\Omega\setminus\overline{V}$. These sets are chosen such that $V$ intersects $\Gamma$ and that $\partial(\Omega\setminus\overline{V_{j}})$ are Lipschitz continuous.
We start by proving that $R(L_{V_{1}}(A))$ and $R(L_{V_{2}}(A))$ have a trivial intersection. Hence, let $\varphi\in R(L_{V_{1}}(A))\cap R(L_{V_{2}}(A))$, i.e. $\varphi=w_{j}|_{\Gamma}$ with $w_{j}=w_{F_{j}}^{A}$ for some $F_{j}\in L^{2}(V_{j})^{d}$ for $j\in\{1,2\}$. From (6.1) and the weak form of the conductivity problem (2.2), with conductivity $A^{*}$, we thus have
@@
$\nu\cdot(A^{*}\nabla w_{j})$ $=0\text{ on }(\partial\Omega)\setminus\overline{V_{j}}.$
@@
The following result shows that the sequence $(f_{j})$ of Neumann conditions (e.g. from Theorem 6.3) such that the corresponding power $\lvert\nabla u_{j}\rvert^{2}$ is localized, can be chosen independently of the function values of $A$ on the set where the power tends to zero.
@@
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset U\cap D$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
@@
$\Lambda^{-}_{C}\geq\Lambda(A_{D})\geq\Lambda^{+}_{C}\quad\text{\emph{implies}}\quad D\subseteq C.$
Case 1
Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{+}_{C}}$ satisfy
@@
Case 2
Using Theorem 6.3 and Proposition 6.4 (as $C\subset\Omega\setminus U$), there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A^{-}_{C}}$ satisfy
@@
Now assume that $D\not\subseteq C$, i.e. that $D^{\bullet}\not\subseteq C$. Since both sets are closures of opens sets and have connected complements, there is a relatively open set that connects $D^{\bullet}\setminus C$ with $\Gamma$ that we may use for localization. Specifically, there exist a relatively open connected set $U\subset\overline{\Omega}\setminus C$ that intersects $\Gamma$ and an open ball $B\subset U\cap D$. Moreover, Assumption 3.4(iii) ensures that we can pick $B$ and $U$ such that we arrive at one of two cases:
@@
$\textup{D}\mkern-1.5mu\Lambda^{-}_{C}\geq\Lambda(A_{D})-\Lambda(A_{0})\geq\textup{D}\mkern-1.5mu\Lambda^{+}_{C}\quad\text{\emph{implies}}\quad D\subseteq C.$
@@
Case 1
We now combine Proposition 5.5 with Proposition 5.6(ii), using the facts that $A_{D}-A_{0}$ is positive semidefinite in $U$ and $A_{D}-A_{0}\geq cI$ in $B$ for some scalar $c>0$, which results in
@@
Case 2
We use Proposition 5.5 together with the facts that $A_{D}-A_{0}$ is negative semidefinite in $U$ and $A_{D}-A_{0}\leq-cI$ in $B$ for some scalar $c>0$, which results in
@@
Next, assume $B\not\subset D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy
@@
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{\emph{implies}}\quad B\subset D^{\bullet}.$
@@
Next, assume $B\not\subset D^{\bullet}$. Since $D^{\bullet}$ is closed and has connected complement, while $B$ is a non-empty open set, there exist a relatively open and connected set $U\subseteq\overline{\Omega}\setminus D^{\bullet}$ intersecting $\Gamma$ and $B$, and an open ball $\widehat{B}\subset U\cap B$. From Theorem 6.3 there is a sequence $(f_{j})$ in $L^{2}_{\diamond}(\Gamma)$ such that $u_{j}=u_{f_{j}}^{A_{0}}$ satisfy